NASA International Space Station On-Orbit Status 10 November 2016

©NASA

NASA International Space Station On-Orbit Status 10 November 2016.

The three Expedition 50 crew members today are working to unload cargo and set up new computer servers onboard the station. In Kazakhstan, the next trio to launch to the International Space are continuing their final prelaunch preparations.

Commander Shane Kimbrough completed work on Wednesday on the Oxygen Generation System (OGS) in the U.S. side of the International Space Station. The OGS was down for maintenance as he and ground specialists troubleshot the device due to a low voltage signature. Teams in Mission control conducted checkouts of the system today and reactivated it successfully, enabling OGS to once more supply oxygen to the space station crew. Final checkouts on the system are on the plan for Friday.

Today, Kimbrough set up new laptop computers to allow ground specialist to load required software. He also unloaded additional cargo from the attached Orbital ATK Cygnus cargo craft.

Cosmonauts Sergey Ryzhikov and Andrey Borisenko performed routine maintenance on the ventilation system inside the Russian segment and took part in the Pilot-T experiment, which assesses the cosmonauts ability to perform complex tasks at different points during their spaceflight.

Three new station crew members are in Kazakhstan preparing for a Nov. 17 launch to the station. Veteran station residents Peggy Whitson of NASA and Oleg Novitskiy of Roscosmos, along with first-time space flyer Thomas Pesquet of the European Space Agency will take a two-day, 34-orbit trip to the station after their liftoff and stay in space till May. The group reviewed flight plans and took part in traditional preflight events, including a media day and the ceremonial raising of flags and planting of trees.

On-Orbit Status Report

Oxygen Generation System (OGS) Maintenance: This morning the crew took a return-to-ground sample of the OGS Recirculation Loop and restowed all tools and hardware the crew has used the last 2 days for the OGS maintenance. The calibration settings for the OGA were updated based on performance this morning and the unit is now in Process, producing 2 kg of O2 per day. The current plan is to stay in this configuration overnight. Tomorrow, OGA will be put in a safe configuration to allow the crew to inspect the H2 ORU QD.

Glacier-1 Buffer Troubleshooting: On October 17, 2016 Glacier-1 internal memory disk buffer reached 99% and was not responding to clean up commands. Glacier-1 was unpowered to prevent the unit from reaching max capacity on the memory disk and risk file corruption. Today, the crew connected the Glacier into the EXpedite PRocessing of Experiments to Space Station (EXPRESS) Rack-2 Laptop Computer (ELC) for ground controllers to command through Remote Desktop Protocol (RDP) and successfully cleared the internal memory disk buffer to below 70%. Glacier-1 is currently not in use and the next planned science operations are scheduled for SpX-10.

Fluid Science Laboratory (FSL) Software Update: To overcome the non-functioning high rate data link nominally used to downlink data from the FSL, the crew connected the Multi-purpose Computer and Communications (MPCC) Laptop to the FSL via S-Video cable. The crew was unsuccessful in updating the software on the MPCC laptop, possibly due to a file error. Ground teams are evaluating the software files used during the update. European Space Agency (ESA) plans to have the Simulation of Geophysical Fluid Flow Under Microgravity-2 (Geoflow-2) experiment's video and data downlink through the Joint Station Local Area Network (JSL) by connecting it to the MPCC. The Geoflow-2 experiment, first run is scheduled for November 14. Geoflow-2 studies heat and fluid flow currents within the Earth's mantle. Geoflow-2 aims to improve computational methods that scientists and engineers use to understand and predict the processes in the Earth's mantle that lead to volcanic eruptions, plate tectonics and earthquakes.

Station Support Computer Server Load: The crew deployed and configured the new SSC Server2 ZBook laptop to allow ground specialist to load required software. The ground completed the SSC Server2 load. This afternoon, the crew moved the Ethernet and power cables to SSC Server1 and prepared the laptop for ground teams to load that laptop. SSC Server1 will host ISS-Server1 after the transition planned on Friday. SSC Server1 and 2 are the first use of the new ZBook model laptop onboard ISS.

Orbital ATK (OA)-5 Cargo Operations: The crew performed 1.5 hours of Cygnus cargo transfer operations today. As of the last report the crew had completed 19.5 hours of cargo transfer with an estimated 7.5 hours remaining.

International Docking Adapter (IDA) Survey: The robotics team completed the dynamic surveys that simulate the view of an incoming docking vehicle as it approaches the recently installed IDA at Pressurized Mating Adapter 2 (PMA-2). These surveys used the Special Purpose Dexterous Manipulator (SPDM) Camera Light Pan Tilt Assembly (CLPA) and pre-configured automatic trajectories to simulate vehicle motion. The approaches were performed in various lighting conditions.

Today's Planned Activities
All activities were completed unless otherwise noted.

In Flight Maintainence (IFM) Oxygen Generator System (OGS) Sample
Soyuz 732 Samsung Tablet Recharge, Initiate
SM Ventilation Subsystem Preventive Maintenance. Group В2
Finding the cause of current overload in FGB ПШО31 exchange bus via SM-FGB Docking Assembly -X.
Oxygen Generation System (OGS) Photodocument and Restow
GLACIER Cable Connect
FSL MPCC Laptop Installation Hardware Pre-Gathering
FSL Stowage Container Opening and S-Video Cable Retrieval
FSL Stowage Container Closure
MPCC1 USB Stick move from SSC11 Laptop to MPCC Laptop
FSL MPCC Laptop Relocation, Powering On & Software Image Installation
PILOT-T. Experiment Ops.
Station Support Computer Server 2 Load Prep
Soyuz 732 Samsung Tablet Recharge, Terminate
Move Cables between Station Support Computer Servers
Station Support Computer Server 2 Load Prep
СОЖ maintenance
Progress 433 (DC1) Transfers and IMS Ops
Crew Handover Conference
GLACIER Cable Disconnect
Transfer Cygnus Cargo Operations
Cygnus Cargo Operations Conference
FSL MPCC Laptop Reboot
FSL MPCC Laptop Hardware Relocation and Powering on Laptop
FSL AVM brackets installation & Facility Core Element release in preparation of scientific operations
Station Support Computer Server 1 Tear Down
IMS Update

Completed Task List Items
None

Ground Activities
All activities were completed unless otherwise noted.
4B3 Battery Reconditioning
SSC Server1 and 2 software load
OGS activation

Three-Day Look Ahead:
Friday, 11/11: GCM Troubleshooting, PFS PuFF Calibration Syringe Maintenance
Saturday, 11/12: Weekly Housekeeping, Crew Off Duty
Sunday, 11/13: Crew Off Duty

QUICK ISS Status - Environmental Control Group:
Component - Status
Elektron - On
Vozdukh - Manual
[СКВ] 1 - SM Air Conditioner System ("SKV1") - Off
[СКВ] 2 - SM Air Conditioner System ("SKV2") - Off
Carbon Dioxide Removal Assembly (CDRA) Lab - Standby
Carbon Dioxide Removal Assembly (CDRA) Node 3 - Operate
Major Constituent Analyzer (MCA) Lab - Operate
Major Constituent Analyzer (MCA) Node 3 - Idle
Oxygen Generation Assembly (OGA) - Process
Urine Processing Assembly (UPA) - Standby
Trace Contaminant Control System (TCCS) Lab - Off
Trace Contaminant Control System (TCCS) Node 3 - Full Up

Please follow SpaceRef on Twitter and Like us on Facebook.